OpenStudy (anonymous):

∠ RQS and ∠ TQP are vertical angles where m∠ RQS = 3(x + 7) and m∠ TQP = 4x + 15. 1. Solve for x. 2. Find m∠ RQS and m∠ TQP 3. Show how you can check your answer.

OpenStudy (anonymous):

Do I start this by combining like terms?

jimthompson5910 (jim_thompson5910):

∠ RQS and ∠ TQP are vertical angles so this means that the two angles are congruent

jimthompson5910 (jim_thompson5910):

so... m∠ RQS = m∠ TQP 3(x + 7) = 4x + 15 x = ???

OpenStudy (anonymous):

3x + 21 = 4x + 15

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

3x + 21 = 4x + 15 - 21 -21 = 3x = 4x - 6

OpenStudy (anonymous):

3x/3 = 4x/3 - 6 x = -6

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

actually sorry

jimthompson5910 (jim_thompson5910):

it should be -x = -6 so x = 6

jimthompson5910 (jim_thompson5910):

you were close though

OpenStudy (anonymous):

oh okay well then I plug that into the equations right?

jimthompson5910 (jim_thompson5910):

yes into each angle equation

OpenStudy (anonymous):

Thanks :)

jimthompson5910 (jim_thompson5910):

sure thing